OpenStudy (anonymous):

If f(x)=x^2+3x+5 , what is f(3+h) ? Please help

OpenStudy (anonymous):

@tom982 do you think you could help me with this one ?

OpenStudy (anonymous):

oh hell naahhhh

OpenStudy (anonymous):

Lol ahh I know it's insane

OpenStudy (anonymous):

All you have to do is replace the x with (3+h): f(x)=x^2+3x+5 f(3+h)=(3+h)^2+3(3+h)+5 =9+6h+h^2+9+3h+5 =h^2+9h+23